OpenStudy (anonymous):

I need to find the perimeter of an Annulus. The radius of the large circle is 2, small circle is 1. The area of the annulus is 9.4 square units. Anyone know how to find out the perimeter? :(

OpenStudy (anonymous):

An annulus is simply a ring - you need to add the circumference of the outside circle and the inside circle (the "hole" in the middle of the ring.) \[Total C=2\pi{r_1}+2\pi{r_2}\] \[=2\pi{({r_1+r_2})}\] <-- \[{r_1}\] is the outer radius. Plug and chug.